Fermi National Laboratory
 

Accelerator Update for May 18, 2009 - May 20, 2009

 

Monday May 18

The day shift began with Operations monitoring the Tevatron (TeV) store 7057 (5.21E31) and the Antiproton Source stack (11.6E10), with the Recycler and Main Injector (MI) off for access to replace a failed MI magnet, with the Linac and Booster off due to MI work, with the Meson MTest, MiniBooNE, and NuMI experiments off due MI work, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Support personnel had the quadrupole magnet at MI-42 installed and its vacuum leak-checked by 9 AM.

Operations terminated store 7057 at 9 AM. A D0 access followed and continued through out the day shift.

RF personnel reported that they had swapped out the bad power amplifier for a RF station (MIRF13). Also, technicians filled the kicker with fluorinert cooling fluids.

MI was out of access around noon.

Operations established a MI ramp around 2 PM.

Operations established beam to MiniBooNE by 2:15 PM.

Operations had stable beam established to Pbar, NuMI, MiniBooNE, and MTest by 3 PM. Operators monitored the vacuum at MI-42 and increased beam intensities as appropriate.

The D0 controlled access was completed by 8 PM.

Operations disabled the MTest beam at 9 PM.

Operations put the TeV into HEP shot setup at 11:15 PM.

Tuesday May 19

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (181.2E10) and the stack (34.6E10), and with NuMI and MiniBooNE taking beam.

Operations established store 7061 by 12:55 AM with an average initial luminosity of 20.3E31.

Operations established beam to MTest by 2:15 AM.

MI personnel conducted coalescing studies all afternoon.

Operations terminated store 7061 at 5 PM and then put the TeV into HEP shot setup. Operations also disabled the MTest beam.

The TeV quenched during shot setup. Operators and TeV personnel investigated for the rest of the evening shift.

A Recycler stash of 100E10 was lost at 9:15 PM. A cause could not be determined for the sudden beam loss. As a check, Proton were injected into Recycler for an orbit check; also an antiproton transfer was injected into the Recycler. There were no problems.

The TeV cryo system had recovered by 11:40 PM.

Wednesday May 20

The midnight shift began with the TeV quench investigation continuing, with Operations monitoring the stash (62.0E10), with Pbar stacking (14.3E10), and with NuMI and MiniBooNE taking beam.

TeV personnel made a horizontal change at TeV sector F0, lowered the chromaticities in the squeeze sequence, and adjusted the D0 colliding beam crossing angle.

Operations established beam to MTest at 4:20 AM.

Operations put the TeV into HEP shot setup at 4:40 AM.

Operations established store 7066 at 6:15 AM with an average initial luminosity of 16.9E31.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/20/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory